# Birgir Hrafnkelsson Morris Riedel Egill Skúlason

### v1

@@ 4.3.2 Hybrid Climbing Image NEB with Minimum Mode Following (CI-NEB-MMF)
While the standard Climbing Image NEB (CI-NEB) method is effective at converging to a saddle point, the final stages of relaxation for the climbing image can be slow, particularly on flat potential energy surfaces. To accelerate this final convergence, a hybrid approach has been implemented that integrates a dedicated minimum-mode following (MMF) saddle search directly into the NEB optimization cycle. This method can be seen as a two-stage refinement strategy within each NEB iteration once the path is sufficiently relaxed 151515gh-230 to EON.
@@ 4.3.3 Case Study: Isomerization of Ethylene Oxide to Acetaldehyde
The reaction involves the rearrangement of ethylene oxide, a three-membered cyclic ether (epoxide), into its more stable isomer, acetaldehyde. The primary thermodynamic driving force for this exothermic reaction is the release of significant ring strain present in the ethylene oxide molecule. The C-C-O bond angles in the epoxide ring are constrained to approximately 60°, a severe deviation from the ideal 109.5° for sp³-hybridized atoms. This strain makes ethylene oxide a high-energy, reactive species. The rearrangement allows the ring to open, forming the more stable carbonyl and methyl groups of acetaldehyde and releasing the stored strain energy.
@@
Figure 11: Comparison of the optimization process for the ethylene oxide to acetaldehyde isomerization using (a) the hybrid CI-NEB-MMF and (b) the standard CI-NEB methods. Each colored line represents the reaction path at a specific point in the optimization, progressing towards the final, converged path (based on the climbing image) shown in black. The reaction coordinate on the x-axis is defined as the cumulative Cartesian distance (Å) between successive images along the path. While both methods find the identical transition state estimate, the color bars highlight the significantly greater efficiency of the hybrid method, which converges in approximately 70 steps, whereas the standard method requires over 250 steps.
Figure 12: 2D landscape projection of the converged hybrid CI-NEB-MMF path. This visualization, developed for this work, plots the trajectory on a coordinate system of RMSD from the reactant vs. RMSD from the product. The interpolated energy contours reveal the topography of the potential energy surface traversed by the path.
@@
The reaction was modeled using the PET-MAD [50] machine learning potential (v1.1.0) through the novel Metatomic interface [78] implemented in EON. The entire process, from initial path generation using the IDPP [31] method to the final NEB calculations, was automated using a Snakemake workflow [87]. Both the standard and hybrid NEB calculations started from identical, pre-minimized endpoints and an IDPP-generated [31] initial path from ASE [74]. An L-BFGS optimizer was used to relax the path until the maximum force on any image fell below the convergence criterion of 0.01 eV/Å. In both calculations, the climbing image was activated once the path was sufficiently relaxed. For the hybrid method, the local minimum-mode following refinement was also triggered at this point, with up to 10 minimum mode following (MMF) steps applied to the climbing image during each subsequent NEB iteration.
@@
Standard CI-NEB 371 7444
Hybrid CI-NEB-MMF 68 2754
The hybrid approach required approximately 63% fewer Potential Energy Surface calls and converged in 82% fewer NEB steps. This efficiency gain stems from the MMF method’s ability to rapidly converge the climbing image to the saddle point once the NEB path is in the correct region, avoiding the slow relaxation characteristic of standard CI-NEB on flat or nearly flat potential energy surfaces.
@@ 4.4 Workflow engines
Another critical function of a workflow engine in a scientific context is the ability to programmatically encode best practices and enforce reproducibility. Many computational methods, such as the NEB technique, require a specific sequence of preparatory and execution steps for reliable results. Manually performing this sequence can be tedious and prone to human error, such as forgetting a critical step. By defining the entire protocol as a series of dependent rules, Snakemake transforms a manual checklist into a robust, automated, and self-documenting scientific component.
For example, a best-practice NEB workflow, as illustrated in Figure 13, involves several distinct stages: first, ensuring the reactant and product structures are fully minimized; second, generating a sensible initial path between these endpoints; and only then, executing the main NEB optimization. By encoding this logic in a Snakefile, one can guarantee that the initial path is never generated with unrelaxed endpoints, thus preventing erroneous calculations and embedding expert knowledge directly into the computational tool. This approach ensures that every calculation is performed consistently and correctly, forming the foundation of truly reproducible research.
@@ 5 Efficient Gaussian Process Regression
This chapter presents the gpr_optim as a concrete first step towards solving this workflow problem for scientific discovery. The core contribution of this work is afterall, not only to achieve state of the art 161616always a moving target performance in terms of the number of calculations, but also to provide an architectural blueprint that demonstrates how to move away from monolithic, single-purpose applications towards an interoperable ecosystem.
@@ 5.1 Design
Finally the dimer itself is translated and rotated through the L-BFGS, which is indpendently implemented within the codebase.
@@ 5.2 Surface systems
To handle extended systems, a finite cutoff is taken for determining pairs. Active pairs are updated on each new iteration. For molecular systems, the cutoff is arbitrarily large. The copper hydrogen dissociation in Figure 17 has been the unit test for the development, and the MATLAB results of 230.6 seconds dropped reliably to 12.9 seconds  [^17] From gproptim at commit a4d1fdaa0ed943d0c9e8b2931db12a4148be0ba4.
@@
While effective, the common approach of using a simple radial cutoff to define the local environment can fail in realistic systems [83]. Structural defects, interfaces, or thermal fluctuations cause the number of atoms within a fixed radius to vary, creating an incompatibility with machine learning models that require a fixed-size input vector.
To resolve this, the radial cutoff has been replaced with a nearest-neighbor selection scheme [^18] From HaoZeke/gproptim at a837ec75f537c551d058e7a170ec880031879f52. This approach guarantees a fixed number of atoms are chosen to represent the local environment, ensuring a consistent number of degrees of freedom for the model’s input. The environment is constructed by selecting a constant number of atoms that are closest to any of the primary moving atoms (e.g., the dissociating molecule or a hopping adatom).
The choice of the number of neighbors is a critical hyperparameter that directly controls the computational scaling of the method. However, sensitivity analyses show that simply increasing the number of atoms does not necessarily improve model accuracy, making a careful and parsimonious selection essential for both efficiency and predictive power.
5.3 Data dredging
@@
It is also important to note that MMF methods are often “finishing” methods, after a reasonable guess to the saddle has been generated. To this end, many single ended searches in practice are rarely over an eV or two away in energy from the nearest minima or point of interest. We show in Figure 18 that the distribution of energy differences $(\Delta E)$ is excessively broad, ranging over
Figure 18: Distribution of Energetic and Structural Properties for the Sella Transition State Dataset. (Top) Probability density of energy differences: $\Delta$(S, I) between the final saddle and initial geometry, $\Delta$(S, M) between the final saddle and the minimized initial geometry, and $\Delta$(I, M) between the initial geometry and its minimized form. The distributions, particularly for $\Delta$(S, I), are extremely broad, spanning nearly 20 eV. This range far exceeds physically realistic energy barriers for the included reaction types, suggesting many initial geometries are highly unstable and unrepresentative of approximate saddle points. (Bottom) Probability density of the root-mean-square deviation (RMSD) between the initial (I) and final saddle (S) structures. The distribution is highly skewed, with a median of 0.36 Å and a significant tail extending to large structural deviations. This, coupled with the wide energy distributions, indicates that the optimization process often involves large, chemically questionable geometric changes rather than the refinement of a reasonable guess. Plotted from data published in [36].
@@ 5.4 Performance characteristics
A comparative analysis of the GPDimer and Sella methods was conducted on a subset of 345 systems for which both optimizers converged to the same saddle point, defined as having an energy difference of less than 0.01 eV. The primary metric for comparison is the number of HF calculations required to reach convergence. On average, the performance of the two methods is comparable. The GPDimer method required a median of 29 HF calculations, while Sella required a median of 31. Notably, the GPDimer achieves this efficiency using Cartesian coordinates, whereas Sella employs internal coordinates, which are generally considered more suitable for the varied stiffness of molecular degrees of freedom. The GPDimer proved more efficient in 57% of the cases, on average reducing the computational cost by 8 HF calculations relative to Sella.
Aggregate statistics alone obscure a key relationship. Figure 19 reveals how computational cost depends on the quality of the initial guess. For initial structures close to the final saddle point, possessing a RMSD below 0.6 Å, the GPDimer consistently requires fewer HF calculations. Its efficiency advantage diminishes as the initial structural deviation increases. In the intermediate range, between 0.6 Å and 1.2 Å RMSD, the methods perform almost identically. Here, Sella’s internal coordinates likely aid the more complex optimization. For the most difficult cases, with initial displacements greater than 1.2 Å, the GPDimer again demonstrates superior efficiency. This suggests a greater robustness when converging from a poor initial structure.
Figure 19: Ridgeline plot showing the distribution of Hartree-Fock (HF) calculation counts required for convergence for the GPDimer (blue) and Sella (red) methods. The data fall into bins according to the root-mean-square deviation (RMSD) between the initial and final saddle geometries. At low RMSD values (< 0.6 Å), GPDimer shows clear efficiency. The methods perform comparably in the intermediate RMSD range. At high RMSD values, GPDimer again holds a performance advantage. The visualization confirms that algorithm efficiency depends strongly on the quality of the initial guess. Plotted from data published in [36].
@@ 5.5 Cataloging saddles
As a diagnostic, the NEB (described in Section 2.2 and Section 4.3.2) can be used to determine the quality of the saddles found by sella and the GPDimer. For the first system in the benchmark, singlet 000, a 16-atom acyclic ether ($\mathrm{C}_{5}\mathrm{OH}_{10}$), both sella and GPDimer calculations are launched until convergence. The GPDimer identifies a hydrogen transfer saddle point with a RMSD of just 0.2 Å after only 23 HF calculations, while Sella locates a saddle corresponding to a methyl group rotation which is lower in energy by 0.4 eV, but the configuration is significantly further away (RMSD of 0.6 Å).
@@
To explore the connectivity between the initial geometry and the Sella saddle, we employ the MMF-NEB protocol (Section 4.3.2), leveraging the fact that NEB bands can be formed between arbitrary points. The resulting optimization history, plotted against the 1D path coordinate (defined in Section 4.2, Eq. 49), is shown in Figure 20. The “reactant” configuration is the initial configuration (at $s=0$, $E=0\text{ eV}$), and the “product” is the Sella-located saddle (at $s\approx 5.2\text{ \AA }$, $E\approx-1.3\text{ eV}$).
@@
Figure 20: The optimization history of an NEB connecting the initial reactant (left, $s=0$) to the Sella-located saddle point (right, $s\approx 5.2\text{ \AA }$). The path is fairly baroque, first relaxing barrierlessly into a deep intermediate minimum ($E\approx-5.3\text{ eV}$). It then climbs over a newly-identified transition state (an inflection at $s\approx 2.4\text{ \AA }$) to reach the Sella saddle.
The trajectory reveals a landscape far more complex than a simple ascent. The path does not lead directly to the Sella saddle. Instead, it first follows a steep, barrier-free relaxation from the high-energy reactant ($E=0\text{ eV}$) into a previously uncharacterized, deep intermediate minimum at $s\approx 1.1\text{ \AA }$ ($E\approx-5.3\text{ eV}$). From this newly discovered basin, the path must then climb out over a newly identified transition state (the inflection point at $s\approx 2.4\text{ \AA }$, $E\approx-0.9\text{ eV}$) to finally reach the Sella saddle.
The 2D landscape projection (Figure 21), generated using the methods from Sec. 4.2, confirms this topography. This second NEB calculation maps the MEP specifically for the relevant event: the path from the initial reactant to the deep intermediate minimum. This “top-down” view clearly shows the reactant on a high-energy plateau (yellow contour, $E\approx 0\text{ eV}$) and the minimum in a deep valley (dark teal, $E\approx-5.3\text{ eV}$). The path (colored line) clearly traverses a small energy ridge, the highest point of which is marked by the white star. This star, which coincides exactly with the transition state for this path, is the GPDimer saddle, confirming its identity as the correct, proximal barrier for the initial relaxation.
@@
Figure 21: A 2D landscape projection of the MEP connecting the initial reactant (top-left) to the deep intermediate minimum (bottom-right). The white star explicitly marks the GPDimer saddle, which coincides exactly with the true transition state on this path, separating the initial configuration from the deep minimum.

### v2

@@ 4.3.2 Hybrid Climbing Image NEB with Minimum Mode Following (CI-NEB-MMF)
While the standard Climbing Image NEB (CI-NEB) method is effective at converging to a saddle point, the final stages of relaxation for the climbing image can be slow, particularly on flat potential energy surfaces. To accelerate this final convergence, a hybrid approach has been implemented that integrates a dedicated minimum-mode following (MMF) saddle search directly into the NEB optimization cycle. This method can be seen as a two-stage refinement strategy within each NEB iteration once the path is sufficiently relaxed 171717gh-230 to EON.
@@ 4.3.3 Case Study: Isomerization of Ethylene Oxide to Acetaldehyde
The reaction involves the rearrangement of ethylene oxide, a three-membered cyclic ether (epoxide), into its more stable isomer, acetaldehyde. The primary thermodynamic driving force for this exothermic reaction is the release of significant ring strain present in the ethylene oxide molecule. The C-C-O bond angles in the epoxide ring are constrained to approximately $60\deg$, a severe deviation from the ideal $109.5\deg$ for \$sp3\$-hybridized atoms. This strain makes ethylene oxide a high-energy, reactive species. The rearrangement allows the ring to open, forming the more stable carbonyl and methyl groups of acetaldehyde and releasing the stored strain energy.
@@
Figure 12: Comparison of the optimization process for the ethylene oxide to acetaldehyde isomerization using (a) the hybrid CI-NEB-MMF and (b) the standard CI-NEB methods. Each colored line represents the reaction path at a specific point in the optimization, progressing towards the final, converged path (based on the climbing image) shown in black. The reaction coordinate on the x-axis is defined as the cumulative Cartesian distance (Å) between successive images along the path. While both methods find the identical transition state estimate, the color bars highlight the significantly greater efficiency of the hybrid method, which converges in approximately 20 steps, whereas the standard method requires over 70 steps.
Figure 13: 2D landscape projection of the converged hybrid CI-NEB-MMF path for the ethoxy acetal system. The trajectory is plotted on a coordinate system of RMSD from the reactant vs. RMSD from the product. Black points represent the discrete sampling history during optimization. The interpolated energy contours (color scale in eV) reveal the topography of the potential energy surface, highlighting the minimum energy path connecting the metastable reactant to the stable product via a high-energy transition state.
While the 1D projection in Figure 12 is essential for comparing optimization efficiency and visualizing the energy barrier, it fundamentally compresses the multidimensional pathway into a single coordinate. We can visualize the geometric progression of the reaction and the topographic context of the Potential Energy Surface with the landscape projection method developed in section 4.2, shown in Figure 13. This projects the converged path onto a coordinate system defined by the RMSD of the endpoints. The trajectory traces the channel from the reactant basin (top-left, $RMSD_{P}\approx 0.8\text{ \AA }$) to the product basin (bottom-right, $RMSD_{R}\approx 0.8\text{ \AA }$).
@@
The reaction was modeled using the PET-MAD [63] machine learning potential (v1.1.0) through the novel Metatomic interface [116] implemented in EON. The entire process, from initial path generation using the IDPP [32] method to the final NEB calculations, was automated using a Snakemake workflow [123]. Both the standard and hybrid NEB calculations started from identical, pre-minimized endpoints and an IDPP-generated [32] initial path from ASE [112]. An L-BFGS optimizer was used to relax the path until the maximum force on any image fell below the convergence criterion of 0.01 eV/Å.
In the standard CI-NEB calculation, the climbing image was activated once the maximum force fell below 0.5 eV/Å. In the hybrid CI-NEB-MMF calculation, the climbing image was activated earlier at 1.5 eV/Å, while the local minimum-mode following refinement was triggered only after the force reached 0.5 eV/Å. Up to 20 minimum mode following (MMF) steps were applied to the climbing image during each subsequent NEB iteration.
@@
Standard CI-NEB 77 1564
Hybrid CI-NEB-MMF 22 582
The hybrid approach required approximately 63% fewer Potential Energy Surface calls and converged in 71% fewer NEB steps. This efficiency gain stems from the MMF method’s ability to rapidly converge the climbing image to the saddle point once the NEB path is in the correct region, avoiding the slow relaxation characteristic of standard CI-NEB on flat or nearly flat potential energy surfaces.
@@ 4.4 Workflow engines
Another critical function of a workflow engine in a scientific context is the ability to programmatically encode best practices and enforce reproducibility. Many computational methods, such as the NEB technique, require a specific sequence of preparatory and execution steps for reliable results. Manually performing this sequence can be tedious and prone to human error, such as forgetting a critical step or using mismatched model versions. By defining the entire protocol as a series of dependent rules, Snakemake transforms a manual checklist into a robust, automated, and self-documenting scientific component.
For example, a best-practice NEB workflow, as illustrated in Figure 14, involves several distinct stages: first, ensuring the reactant and product structures are fully minimized and consistently aligned; second, generating a sensible initial path between these mapped endpoints; and only then, executing the main NEB optimization. Additionally, the workflow enforces reproducibility by automatically fetching and converting the specific version of the machine learning potential required for the simulation. By encoding this logic in a Snakefile, one can guarantee that the initial path is never generated with unrelaxed or unaligned endpoints, thus preventing erroneous calculations and embedding expert knowledge directly into the computational tool. This approach ensures that every calculation is performed consistently and correctly, forming the foundation of truly reproducible research.
@@ 4.5 Towards maximal concurrency
4.6 Conclusions
The algorithmic and architectural developments presented in this chapter serve a specific, unifying purpose: to extend the timescale and complexity of accessible chemical transformations. While standard computational packages offer some solutions for routine equilibrium calculations, the exploration of rare events on high-dimensional potential energy surfaces imposes strict requirements on efficiency, stability, and reproducibility that monolithic software architectures cannot meet.
We have established that the “technical” details of implementation are inseparable from the validity of the physical model. The distinction between geometric and electronic criteria for bonding (Section 4.1.2) enables interpretations of saddle geometries. Similarly, the projection and interpolation methods developed for path visualization (Section 4.2, [122]) provide the necessary topological verification that the computed minimum energy pathways connect the intended basins of attraction, a non-trivial confirmation in complex molecular rearrangements.
Furthermore, the restructuring of the EON framework demonstrates that algorithmic efficiency determines scientific feasibility. We demonstrate this through the hybrid CI-NEB-MMF optimizer (Section 4.3.2), which reduces the computational cost of double ended saddle point searches by an order of magnitude 181818submission in progress to Frontiers chemistry.. Thus, modularization in code allows for the determination of transition states with machine learned interatomic potentials that are otherwise untreatable by standard CI-NEB methods due to pathologies of the landscape or the cost of the electronic structure theory. This concept leads to the more wide ranging concept of a distributed micro-service ecosystem for interopble code, a fom of “chemical BLAS (Basic linear algebra subprograms [125])”, leading to a new epoch of standardization.
Finally, the encoding of these procedures into a formal Directed Acyclic Graph via the Snakemake workflow engine (Section 4.4) moves the methodology beyond manual execution. By strictly defining the dependencies between minimization, alignment, and path optimization, we ensure that the complex simulation protocols are robust against operator error and computationally reproducible.
This infrastructure forms the necessary foundation for the work in the subsequent chapter  [^19] Chronologically developed alongside all chapters.. The application of Gaussian Process Regression to saddle point searches with the GPDimer algorithm requires an underlying engine capable of rapid, asynchronous querying of the potential and robust error recovery. We demonstrated how this software environment minimizes overhead and maximizes modularity while providing diagnostic. We can now move towards addressing the theoretical challenge of constructing data-efficient surrogate models for on-the-fly exploration of the energy landscape, demonstrating state of the art saddle search methods using local Gaussian Process acceleration in a production setting.
@@ 5 Efficient Gaussian Process Regression
This chapter presents the gpr_optim as a concrete first step towards solving this workflow problem for scientific discovery. The core contribution of this work is afterall, not only to achieve state of the art 202020always a moving target performance in terms of the number of calculations, but also to provide an architectural blueprint that demonstrates how to move away from monolithic, single-purpose applications towards an interoperable ecosystem.
@@ 5.1 Design
Finally the dimer itself is translated and rotated through the L-BFGS, which is indpendently implemented within the codebase. The scaling of the Gaussian Process methodology in this chapter proceeds as described in Section 2.4.1, that is, exhibiting $O(M^{2}N^{2})$ complexity for storage and $O(M^{3}N^{3})$ for training dominated by the inversion of the covariance matrix, whereas the inference cost scales linearly with the size of the training set $M$.
@@ 5.2 Surface systems
To handle extended systems, a finite cutoff is taken for determining pairs. Active pairs are updated on each new iteration. For molecular systems, the cutoff is arbitrarily large. The copper hydrogen dissociation in Figure 18 has been the unit test for the development, and the MATLAB results of 230.6 seconds dropped reliably to 12.9 seconds  [^21] From gpr_optim at commit a4d1fdaa0ed943d0c9e8b2931db12a4148be0ba4.
@@
Radial cutoffs are generally used [130] to define the local environment. These can fail in complex systems [131]. Structural defects, interfaces, or thermal fluctuations cause the number of atoms within a fixed radius to vary, creating an incompatibility with machine learning models that require a fixed-size input vector.
To resolve this, the radial cutoff has been replaced with a nvel nearest-neighbor selection scheme [^22] From HaoZeke/gproptim at a837ec75f537c551d058e7a170ec880031879f52. This approach guarantees a fixed number of atoms are chosen to represent the local environment, ensuring a consistent number of degrees of freedom for the model’s input. The environment is constructed by selecting a constant number of atoms that are closest to any of the primary moving atoms (e.g., the dissociating molecule or a hopping adatom).
The number of user defined parameters remains the same, swapping a radial cutoff for the number of neighbors to consider. This directly controls the scaling of the method, though sensitivity analyses show that simply increasing the number of atoms does not necessarily improve model accuracy. Future studies will involve using hyperparameter searches [132] with surrogates on more diverse systems.
The results here highlight the manner in which the nearest-neighbor cutoff effectively decouples the computational cost from the total system size, solving the scalability issue inherent to GPR without succumbing to the instability of radial cutoffs. While the remainder of this thesis focuses on molecular benchmarks, this validation ensures the architecture remains applicable to extended systems for future applications.
5.3 Gas phase molecular benchmark
@@
It is also important to note that MMF methods are often “finishing” methods, after a reasonable guess to the saddle has been generated. To this end, many single ended searches in practice are rarely over an eV or two away in energy from the nearest minima or point of interest. We show in Figure 19 that the distribution of energy differences $(\Delta E)$ is excessively broad, ranging over a span of nearly 20 eV.
Figure 19: Distribution of Energetic and Structural Properties for the Sella Transition State Dataset. (Top) Probability density of energy differences: $\Delta$(S, I) between the final saddle and initial geometry, $\Delta$(S, M) between the final saddle and the minimized initial geometry, and $\Delta$(I, M) between the initial geometry and its minimized form. The distributions, particularly for $\Delta$(S, I), are extremely broad, spanning nearly 20 eV. This range far exceeds expected energy barriers at reasonable temperatures and pressures for the small organic molecules in this dataset, suggesting many initial geometries are highly unstable and unrepresentative of approximate saddle points. (Bottom) Probability density of the root-mean-square deviation (RMSD) between the initial (I) and final saddle (S) structures. The distribution is highly skewed, with a median of 0.36 Å and a significant tail extending to large structural deviations. This, coupled with the wide energy distributions, indicates that the optimization process often involves large, chemically questionable geometric changes rather than the refinement of a reasonable guess. Plotted from data published in [50].
The bottom panel in Figure 19 displays the RMSD deviation between the initial (I) and final saddle (S) structures. The RMSD distribution peaks sharply, with a median near 0.36 Å, yet a non-negligible tail stretches toward large structural deviations. Thus, while most starting guesses lie structurally close to their target saddle, a notable fraction requires substantial geometric rearrangement.
A direct inspection of the energy distributions reveals significant energetic instability in many initial guesses, a result that contrasts with the low median RMSD. The $\Delta(\text{S,I})=E_{S}-E_{I}$ distribution (top panel, green) does not center near 0 eV. Instead, its principal mode lies near 2.5-2.5 eV and a broad tail extends down to about 10-10 eV. This pattern indicates that initial geometries often occupy much higher electronic energy than their converged saddle points.Comparing $\Delta(\text{I,M})$ (red) and $\Delta(\text{S,M})$ (yellow) further quantifies this instability. The $\Delta(\text{I,M})$ ridge shifts to more positive energy relative to $\Delta(\text{S,M})$. This displacement simply reflects the $E_{I}-E_{S}$ energy difference, reproducing the finding from the $\Delta(\text{S,I})$ plot and confirming that initial points frequently have higher energies than the corresponding saddles.
This disconnect between structural proximity and energetic instability introduces a substantive complication. A starting geometry that sits energetically far from the saddle region may not lie in the quadratic basin that leads to that saddle despite having proximal configurations in space (w.r.t the RMSD); instead, the high potential energy can place the start in a region of the potential energy surface that grants access to multiple competing downhill pathways. A local finishing method that follows the local gradient therefore cannot guarantee convergence toward the structurally proximal saddle. The elevated starting energy can allow the optimization to descend along an alternative reaction coordinate and converge to an unrelated minimum or a different stationary point. The tail in the RMSD(I,S) distribution corroborates this concern, as many cases require large structural changes before convergence. In sum, the dataset displays a pronounced disconnect between structural and energetic closeness. Many initial geometries function as highly unstable configurations that happen to lie near a saddle in Cartesian space but fail to occupy the correct energetic basin, which in turn makes them unreliable “reasonable initial structures”.
@@ 5.4 Performance characteristics
A comparative analysis of the GPDimer and Sella methods was conducted on a subset of 345 systems for which both optimizers converged to the same saddle point, defined as having an energy difference of less than 0.01 eV. Although this quantification does not rigorously rule out that the geometries of the saddle point configurations are comparable, the distributional analysis and general conclusions do not change. The primary metric for comparison is the number of HF calculations required to reach convergence. On average, the performance of the two methods is comparable. The GPDimer method required a median of 29 HF calculations, while Sella required a median of 31. Notably, the GPDimer achieves this efficiency using Cartesian coordinates [50], whereas Sella employs internal coordinates, which are generally considered more suitable for the varied stiffness of molecular degrees of freedom [134, 135, 136, 133]. The GPDimer proved more efficient in 57% of the cases, on average reducing the computational cost by 8 HF calculations relative to Sella.
Aggregate statistics alone obscure a key relationship. Figure 20 reveals how computational cost depends on the quality of the initial guess. For initial structures close to the final saddle point, possessing a RMSD below 0.6 Å, the GPDimer often requires slightly fewer HF calculations, as the initial surfaces constructed are not optimal. Its efficiency advantage increases as the initial structural deviation increases. In the intermediate range, between 0.6 Å and 1.2 Å RMSD, the methods perform almost identically. For the most difficult cases, with initial displacements greater than 1.2 Å, the GPDimer again demonstrates superior efficiency.
Figure 20: Ridgeline plot showing the distribution of Hartree-Fock (HF) calculation counts required for convergence for the GPDimer (blue) and Sella (red) methods. The data fall into bins according to the root-mean-square deviation (RMSD) between the initial and final saddle geometries. At higher RMSD values (> 0.6 Å), GPDimer shows clear efficiency. The methods perform comparably in the intermediate RMSD range. At high RMSD values, GPDimer again holds a performance advantage. The visualization confirms that algorithm efficiency depends strongly on the quality of the initial guess. Plotted from data published in [50].
@@ 5.5 Cataloging saddles
As a diagnostic, the NEB (described in Section 2.2 and Section 4.3.2) can be used to determine the quality of the saddles found by sella and the GPDimer. We examine the first system in the benchmark, singlet 000, a 16-atom acyclic ether ($\mathrm{C}_{5}\mathrm{OH}_{10}$), for which both sella and GPDimer converge to distinct saddle points.
The GPDimer identifies a hydrogen transfer saddle point with a RMSD of just 0.2 Å after only 23 HF calculations, while Sella locates a saddle corresponding to a methyl group rotation which is lower in energy by 0.4 eV, but the configuration is significantly further away (RMSD of 0.6 Å).
@@
Since energy alone cannot be considered sufficient for cataloging saddles [137] we explore the connectivity between the initial geometry and the Sella saddle, we employ the MMF-NEB protocol (Section 4.3.2), leveraging the fact that NEB bands can be formed between arbitrary points. The resulting optimization history, plotted against the 1D path coordinate (defined in Section 4.2, Eq. 50), is shown in Figure 21. The “reactant” configuration is the initial configuration (at $s=0$, $E=0\text{ eV}$), and the “product” is the Sella-located saddle (at $s\approx 5.2\text{ \AA }$, $E\approx-1.3\text{ eV}$).
@@
Figure 21: The optimization history of an NEB connecting the initial reactant (left, $s=0$) to the Sella-located saddle point (right, $s\approx 5.2\text{ \AA }$). The pathway reveals that the initial structure is not metastable; it relaxes barrierlessly into a deep intermediate minimum ($s\approx 1.1\text{ \AA }$, $E\approx-5.3\text{ eV}$). To reach the Sella saddle, the system must subsequently climb out of this minimum, crossing a distinct transition state (peak at $s\approx 2.4\text{ \AA }$) before proceeding to the distal Sella endpoint.
The 1D reaction coordinate profile (Figure 21) shows that the “initial” configuration ($s=0$) is highly unstable and relaxes barrierlessly into a deep intermediate minimum ($E\approx-5.3\text{ eV}$). From this localized basin, the system must cross a significant barrier (the peak at $s\approx 2.4\text{ \AA }$) to access the region containing the Sella saddle.The critical finding is revealed when we project these states onto the 2D landscape [122] (Figure 22). The saddle point found by GPDimer (marked by the white star) coincides exactly with the barrier peak at $s\approx 2.4\text{ \AA }$. This indicates that the Cartesian-based GPDimer successfully identified the proximal barrier governing the immediate relaxation of the system, whereas the internal coordinate search bypassed this feature in favor of a distal stationary point. For automated kinetic network discovery (AKMC) as implemented in EON, identifying the immediate connectivity is paramount; skipping proximal barriers in favor of lower-energy distal states leads to disconnected reaction networks and missing kinetic links. Thus, GPDimer’s ability to localize the geometrically nearest saddle ensures the correct step-wise mapping of the potential energy surface. The result also underscores the necessity of characterizing the local gradient manifold when sampling from high-energy initial configurations.
@@
Figure 22: A 2D landscape projection of the NEB optimization trajectory connecting the initial reactant (top-left) to the deep intermediate minimum (bottom-right). The white star explicitly marks the GPDimer saddle, which coincides exactly with the true transition state on this path, separating the deep intermediate minimum from the subsequent product channel. Sella’s trajectory (red/pink path endpoints) overshoots this state to find a lower-energy, but distal, saddle point.
5.6 Conclusions
This chapter presented gpr_optim, an implementation of Gaussian Process Regression designed to accelerate minimum mode following methods. By coupling the dimer method with a local surrogate model, we achieved an order-of-magnitude reduction in the number of electronic structure calculations required to locate transition states compared to standard dimer methods.
A critical finding of this work is that the GPR-Dimer, operating in simple Cartesian coordinates, achieves computational parity with—and often outperforms—state-of-the-art methods utilizing complex internal coordinates (Sella). This is a significant architectural advantage. For surface science and catalysis, defining non-redundant internal coordinates is fraught with difficulties due to coordinate singularities and the need for “ghost atoms.” The results on the molecular benchmark demonstrate that the learning capability of the GPR can effectively compensate for the lack of coordinate sophistication. The model implicitly learns the local Hessian structure that internal coordinates attempt to explicitly encode.
Furthermore, the detailed connectivity analysis using NEB diagnostics revealed that “convergence” is an insufficient metric for success. In complex landscapes, algorithms may converge to physically valid but chemically irrelevant saddle points. The GPR-Dimer showed a robust tendency to locate the proximal saddle point governing the immediate exit from the reactant basin, whereas aggressive optimization in internal coordinates occasionally led to landscape traversal and unrelated isomerizations.
Ultimately, this framework provides a robust, coordinate-system-agnostic engine for saddle point searches. While the cubic scaling of exact GPR remains a theoretical bottleneck, the implementation of local active sets (as demonstrated with the CuH2 unit test) ensures the method remains viable for extended systems.